Daniel Mathews a footballer who is based in Sweden received a fine of $2 500 and a two-match suspension for use of performance-enhancing drugs. Daniel receives a match fee of $5 000 per match from the Sweden Footballers Board, plus additional appearance money of $10 000 per match from Adidas. He also pays $500 per match for comprehensive insurance against injury. In opportunity cost terms, the cost of this incident for Daniel Mathews is:

1
Expert's answer
2021-11-30T17:42:58-0500

"CT=2500+2\cdot(5000+10000-500)=\text {} 31500."